A swimming pool can be filled in 15 hours if water enters through a pipe alone, or in 22 hours if water enters through a hose alone. If water is entering through both the pipe and the hose, how long will it take to fill the pool 2/5 full?
Answer by Alan3354(69443)

A swimming pool can be filled in 15 hours if water enters through a pipe alone, or in 22 hours if water enters through a hose alone. If water is entering through both the pipe and the hose, how long will it take to fill the pool 2/5 full?
------------
The pipe fills 1/15 of the pool per hour.
The hose fills 1/22 per hour.
---
Together, they fill 1/15 + 1/22 = 37/330 per hour
--> 330/37 to fill the pool
To fill it 2/5, = (2/5)*(330/37) hours
= 132/37 hours